OpenStudy (anonymous):

PLEASE HELP WILL GIVE A MEDAL Find the derivative of f(x) = -7/x at x = -3.

OpenStudy (anonymous):

\((\dfrac{-7}{x})'=?\)

OpenStudy (anonymous):

the derivative is 7/x^2

OpenStudy (anonymous):

so replace -3 you get 7/9

OpenStudy (michele_laino):

please apply this rule: \[\frac{ d }{ dx }\left( \frac{ 1 }{ x } \right)=-\frac{ 1 }{ x ^{2} }\]